OpenStudy (anonymous):

determine wether the graph of 5y-x=15 and y+5x=10 are parallel

OpenStudy (anonymous):

Parallel lines have the same slope so: 5y-x=15 and y+5x=10 5y=x+15 y=-5x+10 y=x/5+3 Therefore they are not parallel but perpendicular